OpenStudy (lena772):

The label on the car's antifreeze container claims to protect the car between −30°C and 130°C. To convert Celsius temperature to Fahrenheit temperature, the formula is C = 5 over 9 (F − 32). Write and solve the inequality to determine the Fahrenheit temperature range at which this antifreeze protects the car. −30 > 5 over 9 (F − 32) > 130; −22 > F > 266 −30 < 5 over 9 (F − 32) < 130; −22 < F < 266 −30 < 5 over 9 (F − 32); −22 < F 5 over 9 (F − 32) < 130; F < 266

OpenStudy (kira_yamato):

C = 5 over 9 (F − 32) C = 130 => F - 32 = 234 => F = 266 C = -30 => F - 32 = -54 => F = -22 −30 < 5 over 9 (F − 32) < 130; −22 < F < 266